(m+n) 2/3 +(m-n) 2/3 =2a 2/3 if acos 3 ​+3acos.sin 2 =m and asin 3 +3acos 2 .sin=n

(m+n)2/3+(m-n)2/3=2a2/3
if acos3​+3acos.sin2=m
and asin3+3acos2.sin=n

Grade:12th pass

2 Answers

Y RAJYALAKSHMI
9 years ago
m + n = acos3x + 3a cosx sin2x + asin3x + 3a cos2xsinx = a(sinx + cosx)3
(m + n)2/3 = a2/3 (sinx + cosx)2
m – n = acos3x + 3a cosx sin2x –  asin3x –  3a cos2xsinx = a(cosx + sinx)3
 
(m –  n)2/3 = a2/3 (cosx –  sinx)2
(m + n)2/3 + (m –  n)2/3 = a2/3 (sinx + cosx)2 + a2/3 (cosx –  sinx)=
a2/3[(sinx + cosx)(cosx –  sinx)2] = 2a2/3
 
 
gautam
6 years ago
 
 
m + n = acos3x + 3a cosx sin2x + asin3x + 3a cos2xsinx = a(sinx + cosx)3
(m + n)2/3 = a2/3 (sinx + cosx)2
m – n = acos3x + 3a cosx sin2x –  asin3x –  3a cos2xsinx = a(cosx + sinx)3
 
(m –  n)2/3 = a2/3 (cosx –  sinx)2
(m + n)2/3 + (m –  n)2/3 = a2/3 (sinx + cosx)2 + a2/3 (cosx –  sinx)=
a2/3[(sinx + cosx)(cosx –  sinx)2] = 2a2/3